when iron is exposed to oxygen (air), it rusts. how many grams of iron are needed in order to prepare 12.25 grams of Fe2O3?


1
Expert's answer
2022-03-08T17:29:04-0500

4Fe + 3O2 = 2Fe2O3

According to the equation, n (Fe) = 4/2 x n (Fe2O3)

n = m / M

M (Fe2O3) = 159.7 g/mol

M (Fe) = 55.8 g/mol

n (Fe2O3) = m / M = 12.25 / 159.7 = 0.08 mol

n (Fe) = 4/2 x n (Fe2O3) = 4/2 x 0.08 = 0.16 mol

m (Fe) = n x M = 0.16 x 55.8 = 8.9 g